OpenStudy (anonymous):

a triangle with vertices at a (0,0), b(0,4) and c(6,0) is dilated to yield a triangle with vertices at a"(0,0,) , b(0,10) and c"(15,0). the origin is the center of dilation what is the scale factor of the dilation

OpenStudy (campbell_st):

have you plotted the points...?

OpenStudy (anonymous):

no wasnt sure were to start

OpenStudy (campbell_st):



OpenStudy (anonymous):

im not sure what im looking at or for

OpenStudy (campbell_st):

ok... 4 x ? = 10 so ? = 10/4 is that the same as 6 x ?? = 15 so ?? = 15/6 do both equations give the same value... if they do that's the dilation

OpenStudy (anonymous):

oh ok thank you i got it now